### v2

# MCPDepth: Practical Omnidirectional Depth Estimation from Multiple Cylindrical Panoramas via Stereo Matching

Abstract
Omnidirectional depth estimation presents a significant challenge due to the inherent distortions in panoramic images. Despite notable advancements, the impact of projection methods remains underexplored. We introduce Multi-Cylindrical Panoramic Depth Estimation (MCPDepth), a novel two-stage framework designed to enhance omnidirectional depth estimation through stereo matching across multiple cylindrical panoramas. MCPDepth initially performs stereo matching using cylindrical panoramas, followed by a robust fusion of the resulting depth maps from different views. Unlike existing methods that rely on customized kernels to address distortions, MCPDepth utilizes standard network components, facilitating seamless deployment on embedded devices while delivering exceptional performance. To effectively address vertical distortions in cylindrical panoramas, MCPDepth incorporates a circular attention module, significantly expanding the receptive field beyond traditional convolutions. We provide a comprehensive theoretical and experimental analysis of common panoramic projections—spherical, cylindrical, and cubic—demonstrating the superior efficacy of cylindrical projection. Our method improves the mean absolute error (MAE) by 18.8% on the outdoor dataset Deep360 and by 19.9% on the real dataset 3D60. This work offers practical insights for other tasks and real-world applications, establishing a new paradigm in omnidirectional depth estimation. The code is available at https://github.com/Qjizhi/MCPDepth.
1 Introduction
Figure 1: Comparison of stereo images among common panoramic projections.
Depth estimation is a pivotal challenge in geometric computer vision, playing a critical role in 3D scene understanding and robotic perception. Despite substantial advancements achieved through convolutional neural networks (CNNs) in processing perspective images, the task of estimating omnidirectional depth remains particularly challenging due to the severe geometric distortions inherent in panoramic representations. Recent research has investigated both monocular [45, 14] and stereo [47, 18, 19] approaches, each presenting unique advantages and limitations. Methods that apply conventional CNNs to spherical projections [45, 14, 47] often struggle to effectively manage these distortions, while those that directly model spherical epipolar geometry [18] encounter significant computational complexities. Although some strategies have introduced customized convolutional techniques, such as deformable convolution [42], EquiConvs [8], and spherical convolution [19], their practical deployment on resource-constrained robotic platforms remains a formidable challenge [55, 39]. Additionally, the inherent ambiguities associated with single or dual-view depth estimation frequently result in unreliable outputs, further complicating the task.
Several works have explored multi-view approaches, such as SweepNet [53] and OmniMVS [52], which use fish-eye cameras to capture a panoramic field of view (FoV). However, these methods face limitations, including ineffective feature extraction due to severe radial distortions and ultra-wide FoV when using standard 2D convolutions, and incomplete depth reconstruction due to blind spots in fish-eye camera configurations, leading to discontinuities in the spherical cost volume representation.
Recent research has advanced stereo matching for depth prediction by leveraging epipolar constraints to reduce the search space and improve accuracy. Notable contributions include 360SD-Net [47] and MODE [19], which have addressed challenges in complex depth estimation scenarios. MODE, in particular, introduces a two-stage framework that utilizes Cassini projection [50] to simplify epipolar geometry, followed by multi-view depth map fusion to enhance robustness. While these methods achieve state-of-the-art results, they face computational bottlenecks, especially on resource-constrained devices [39], due to their reliance on spherical convolutions [6]. Additionally, Cassini projection introduces significant distortions, particularly near the poles, which can degrade depth map quality.
Despite significant advancements in this field, the influence of projection methods on feature extraction and downstream tasks remains insufficiently explored. As illustrated in Figure 1, different projections exhibit distinct characteristics, each influencing the performance of CNNs and downstream tasks. In this work, we systematically analyze these effects and demonstrate that cylindrical projection is particularly effective for CNN-based feature extraction.
Drawing from the strengths of stereo matching and two-stage frameworks, we propose MCPDepth, a novel framework that leverages cylindrical projection for stereo matching. Our approach offers three key advantages: it significantly reduces geometric distortion compared to spherical projection, it’s compatible with standard 2D convolutions, avoiding computationally intensive spherical convolutions, and it preserves the stereo matching relationship of perspective images, enabling better transfer learning from existing models. Additionally, we introduce a circular attention module that captures long-range dependencies across the full $360{{}^{\circ}}$ vertical field of view (FoV) while mitigating projection-induced distortions. Our contributions can be summarized as follows:
  • We introduce the first framework for omnidirectional depth estimation that leverages stereo matching across multiple cylindrical panoramas.
  • We conduct a comprehensive theoretical and experimental analysis comparing common projections, highlighting the advantages of cylindrical projection.
  • We present an innovative circular attention module designed to alleviate vertical axis distortions in cylindrical panoramas while significantly enhancing the receptive fields of conventional convolutions.
  • Our method sets new benchmarks on the Deep360 (outdoor) and 3D60 (indoor) datasets.
2 Related Work
2.1 Deep Learning-based Stereo Matching
Early methods employed deep neural networks to compute matching costs, such as MCCNN [61], which trains a CNN for initial patch matching costs. Recently, end-to-end neural networks have dominated stereo matching methods. Works such as [28, 30, 22, 10, 57, 21, 41] only use 2D convolutions. Mayer et al. [28] propose the first end-to-end disparity estimation network, DispNet, and its correlation version, DispNetC. Pang et al. [30] introduce a two-stage framework named CRL with multi-scale residual learning. GwcNet [10] proposes the group-wise correlation volume to improve the expressiveness of the cost volume and performance in ambiguous regions.
AANet [57] adopts a novel aggregation algorithm using sparse points and multi-scale interaction. Another series of works [15, 2] use 3D convolutions, which demonstrate great potential in regularizing or filtering the cost volume. GCNet [15] first implements a 3D encoder-decoder architecture aimed at regularizing a 4D concatenation volume. PSMNet [2] proposes a stacked hourglass 3D CNN in conjunction with intermediate supervision to regularize the concatenation volume. Recently, iterative methods [25, 44, 17, 56] have shown impressive results. RAFTStereo [25] proposes to recurrently update the disparity field using local cost values retrieved from the all-pairs correlations. IGEV-Stereo [56] further advances this iterative approach by introducing a geometry encoding volume to encode non-local geometry and context information. Selective-Stereo [49] proposes a novel iterative update operator SRU for iterative stereo matching methods.
In parallel, substantial progress has been made in multi-view stereo (MVS) techniques [59, 58, 3, 9], which focus on generating 3D reconstructions from multiple perspective views, albeit primarily designed for limited-FoV cameras.
2.2 Omnidirectional Depth Estimation
Omnidirectional depth estimation has developed tremendously with neural networks. Zioulis et al. [65] present a learning-based monocular depth estimation method, trained directly on omnidirectional content in the ERP domain, and later propose CoordNet [66] with a spherical disparity model. BiFuse [45] uses both equirectangular and cubemap projections for depth estimation. A more effective fusion framework for ERP and cubemap projection is proposed in Unifuse [14]. Cheng et al. [4] introduce a depth sensing system by combining an OmniCamera with a regular depth sensor. 360SD-Net [47] is the first end-to-end trainable network for stereo depth estimation using spherical panoramas. CSDNet [18] focuses on left-right stereo and uses Mesh CNNs [13] to overcome spherical distortion. SweepNet [53] and OmniMVS [52] use multi-view fish-eye images for omnidirectional depth maps. However, most of them are based on spherical projection and extract spherical features with regular convolutions, and none of them discuss the properties of cylindrical projection.
Cheng et al. [4] propose a spherical feature transform layer to reduce the difficulty of feature learning. MODE [19] adopts spherical convolution from Spherenet [6], but the customized CUDA implementation poses deployment challenges on robotic platforms [39].
Jun et al. [38] employ cylindrical panoramas for stereo matching, but without CNNs or analysis of cylindrical projection properties, and 12 perspective images are stitched to obtain the panoramas.
2.3 Self-Attention Module
Attention mechanisms were first introduced by [1] for the encoder-decoder in a neural sequence-to-sequence model to capture token correspondence between sequences. Self-attention, designed for single contexts, encodes long-range interactions and has been widely applied in computer vision, achieving state-of-the-art performance [43, 31, 48, 62, 12, 29, 32, 40, 5]. Global self-attention in image processing is computationally expensive due to the need to calculate the relationship between every pixel and every other pixel, limiting its practical usage across all layers in a full-attention model. It is shown in [35, 11] that self-attention layers alone could form a fully attentional model by restricting the receptive field of self-attention to a local region.
In stereo matching, CREStereo [17] first adopts the self-attention module from LoFTR [40]. Zhao et al. [64] propose a multi-stage and multi-scale channel-attention transformer to preserve high-frequency information. GOAT [26] uses self-cross attention to capture more representative and distinguishable features. However, these methods are not designed for stereo matching in $360{{}^{\circ}}$ panoramic images.
More recently, some attention mechanisms [24, 37, 60, 63] specifically designed for ERP have been proposed. However, these methods cannot be directly applied to cylindrical projection and face significant deployment challenges and computational overhead.
3 Method
Given $m$ $360{{}^{\circ}}$ cameras, where $m\geq 3$, we have a set of $n={m\choose 2}=\frac{m!}{2!(m-2)!}$ pairs of rectified panoramas $\{(I^{i}_{L},I^{i}_{R})\}^{n}_{i=1}$ with their intrinsic and extrinsic parameters. Our objective is to estimate the omnidirectional depth map $d$ for the left panorama in the first pair $I^{1}_{L}$.
3.1 Preliminaries: Panorama Projections
In this section, we discuss the similarities and differences between spherical and cylindrical projections for stereo matching. Here, cylindrical refers specifically to the vertical cylindrical projection. We omit details of the cubic projection, as it follows the same principles as perspective images. Finally, we provide a comprehensive analysis of the advantages and disadvantages of common panoramic projections.
Figure 2: (a) and (b) compare the spherical and cylindrical projections for stereo matching and their respective epipolar geometries. (c) and (d) represent the schematic drawing of the epipolar plane under spherical and cylindrical projections.
Table 1: Comparison of different projection types. $h$ and $v$ represent the horizontal and vertical FoV, respectively.
Projection Type Advantages Disadvantages
ERP Full coverage: $360{{}^{\circ}}(h)\times 180{{}^{\circ}}(v)$ FoV. Non-linear epipolar geometry, complicating stereo matching.
Cassini Linear epipolar geometry simplifies stereo matching. Full coverage: $360{{}^{\circ}}(v)\times 180{{}^{\circ}}(h)$ FoV. Severe distortion near poles, uneven distortion. Requires custom kernels for processing.
Cubic Linear epipolar geometry. No distortion within individual cube faces. Compatible with standard convolutional kernels. Limited horizontal FoV: $360{{}^{\circ}}(v)\times 90{{}^{\circ}}(h)$. Discontinuities at cube joints, hindering CNN feature learning. Requires fusion module across faces.
Cylindrical Linear epipolar geometry. No distortion along the horizontal axis. Uniform distortion along the vertical axis. Compatible with standard convolutional kernels. Limited horizontal FoV: $360{{}^{\circ}}(v)\times n{{}^{\circ}}(h)$, where $n<180$. Residual distortion along the vertical axis.
As illustrated in Fig. 2, both cylindrical and spherical projections preserve the linear epipolar constraint. In spherical coordinates, $\rho$ represents the Euclidean distance from the origin $O$ to point $P$; $\phi$ is the angle between line $OP$ and the plane $yOz$; and $\theta$ is the angle between line $OP^{\prime}$ and the z-axis, where $P^{\prime}$ is the projection of $P$ on the plane $yOz$. In cylindrical coordinates, $\rho$ denotes the Euclidean distance from the x-axis to point $P$; $\theta$ is the angle between line $OP^{\prime}$ and the z-axis, where $P^{\prime}$ is the projection of $P$ on the plane $yOz$. The conversion between spherical, cylindrical, and Cartesian coordinate systems is illustrated in Eq. 1.
$\begin{aligned} {\color[rgb]{0,0,0}\begin{cases}x=\rho\sin(\phi)\\ y=\rho\cos(\phi)\sin(\theta)\\ z=\rho\cos(\phi)\cos(\theta)\end{cases}}\end{aligned}\begin{aligned} {\color[rgb]{0,0,0}\begin{cases}x=x\\ y=\rho\sin(\theta)\\ z=\rho\cos(\theta)\end{cases}}\end{aligned}$ (1)
The spherical and cylindrical panoramas in Fig. 1 (b) and (f) are generated according to Eq. 2, where $u$ and $v$ are pixel coordinates, $W$ and $H$ are panorama dimensions and $R=H/2\pi$ is the cylinder’s radius, which is the focal length in perspective images. $u$ in the cylindrical panorama is the same as it is in the perspective images.
$\begin{aligned} {\begin{cases}u=(\phi+\frac{\pi}{2})\cdot\frac{W}{\pi}\\ v=(\theta+\pi)\cdot\frac{H}{2\pi}\end{cases}}\end{aligned}\begin{aligned} {\begin{cases}u=-\frac{xR}{\rho}+\frac{W}{2}=-\frac{x}{\rho}\cdot\frac{H}{2\pi}+\frac{W}{2}\\ v=(\theta+\pi)\cdot\frac{H}{2\pi}\end{cases}}\end{aligned}$ (2)
In distortion-free perspective images, an object’s actual length and its pixel length along the horizontal and vertical axes is given by $\Delta u=\frac{f_{x}}{z}\Delta x$ and $\Delta v=\frac{f_{y}}{z}\Delta y$, where $f_{x}$ and $f_{y}$ are the focal lengths along the $x$ and $y$ axes, and $z$ is the distance along the z-axis. Eq. 3 shows these relationships for both spherical and cylindrical projections.
$\begin{aligned} {\begin{cases}\Delta u=f_{\phi}\Delta\phi\approx\frac{f_{\phi}}{\rho\cos{\theta}}\Delta X\\ \Delta v=f_{\theta}\Delta\theta\approx\frac{f_{\theta}}{\rho}\Delta Y\end{cases}}\end{aligned}\begin{aligned} {\begin{cases}\Delta u=\frac{f_{X}}{\rho}\Delta X\\ \Delta v=f_{\theta}\Delta\theta\approx\frac{f_{\theta}}{\rho}\Delta Y\end{cases}}\end{aligned}$ (3)
where $f=R=H/2\pi$. The relation $\Delta u=f\Delta X/\rho$ and approximation $\Delta v\approx f\Delta Y/\rho$ for cylindrical projection hold under the condition that the object is not too large or far from the camera [33]. This approximation means objects in cylindrical projection appear similar regardless of their location. This shift-invariant property facilitates efficient learning by CNNs. In contrast, objects in spherical projection vary with their $\theta$ axis position, limiting the effectiveness of regular convolutions.
In addition, the disparity in spherical projection is defined as angular disparity $d$ (Fig. 2 (c)), where $d=\left|\phi_{l}-\phi_{r}\right|$. This concept has been previously discussed in some works [20, 23, 66, 19]. The relationship between disparity and depth is:
$\rho_{l}=B\cdot\frac{\sin(\phi_{r}+\frac{\pi}{2})}{\sin(d)}=B\cdot\left[\frac{\sin(\phi_{l}+\frac{\pi}{2})}{\tan(d)}-\cos(\phi_{l}+\frac{\pi}{2})\right]$ (4)
where $B$ denotes the baseline. As shown in Fig. 2 (d), the cylindrical projection maintains the same disparity-depth relationship as perspective images:
$\rho_{l}=\frac{B\cdot f}{\left|x_{l}-x_{r}\right|}$ (5)
Tab. 1 summarizes the advantages and disadvantages of different projection types. Cylindrical projection is the most suitable for stereo matching of panoramas due to the following reasons: (1) Compatibility with perspective images: Cylindrical panoramas maintain a disparity definition consistent with perspective images, enabling the direct application of stereo networks originally designed for perspective images. (2) Reduced distortion: Cylindrical panoramas only distort vertically, providing better shift invariance, which enhances CNN feature learning. (3) Simplified Deployment: Spherical panoramas require customized convolutions [42, 6, 8] to extract features. For example, spherical convolutions can’t be exported to widely used ONNX [51] format for deployment, they either need CUDA Plugins for the TensorRT engine on NVIDIA platforms or customized implementation on other embedded devices. Cylindrical panoramas only use regular convolutions, making MCPDepth deployment-friendly.
3.2 Framework
The MCPDepth framework, shown in Fig. 3, includes two stages. In the stereo matching stage, $n$ pairs of rectified cylindrical panoramas (Fig. 3 (a)) are fed into the stereo matching network. The number of pairs ($n$) and cameras ($m$) varies on different datasets: $n=6,m=4$ for Deep360, and $n=3,m=3$ for 3D60. The resulting disparity and confidence maps (Fig. 3 (b)) are reprojected into the Cassini domain with a $180{{}^{\circ}}$ horizontal FoV. The disparity maps are then converted to depth maps. The depth and confidence maps are aligned with the view of $I^{1}_{L}$ using extrinsic parameters as shown in Fig. 3 (c). Black areas indicate invisible and occluded regions.
Figure 3: Framework of MCPDepth. (a) represents 6 pairs of cylindrical panoramas, (b) shows the disparity and confidence maps, and (c) shows the depth and confidence maps. (d) and (e) illustrate the depth map in Cassini and spherical projection.
We use the circular attention module between feature extraction and cost volume with a structure similar to PSMNet [2]. The circular attention module augments the extracted features to capture features from a $360{{}^{\circ}}$ FoV and overcome vertical-axis distortion. These augmented features are then shifted and concatenated to build the cost volume. The disparity map is regressed through the 3D stacked hourglass network. During training, we use the $\ell_{1}$ loss to train the network. The confidence maps are used to measure the reliability of the disparity estimation and are widely used in stereo matching tasks [34]. The confidence map is obtained during inference. Specifically, considering the disparity is obtained through a probability-weighted sum over all disparity hypotheses, we compute the corresponding confidence value by taking a sum of probabilities over the three nearest disparity hypotheses.
We generally follow MODE’s depth fusion stage structure. Specifically, multi-view depth maps, along with their corresponding confidence maps and reference panoramas, are fed into two separate 2D encoder blocks. The fused depth map is then processed through a single decoder block, incorporating skip connections between the encoder and decoder blocks at each scale. The final depth map is generated in the Cassini domain [50] as shown in  Fig. 3 (d), a transverse variant of the equirectangular projection (ERP) commonly used in map projections Fig. 3 (e), but it can be readily converted to the ERP domain. More details are available in the supplementary material.
3.3 Circular Attention
Figure 4: (a) displays the circular attention module in the stereo matching network. (b) represents our attention applied along the vertical axis. $\bigoplus$ denotes element-wise sum. $\bigotimes$ denotes matrix multiplication. Blue boxes are $1\times 1$ convolution and orange boxes are relative positional encoding.
To overcome vertical axis distortion and capture the circular $360{{}^{\circ}}$ features, we introduce a circular attention module. Conventional CNNs have limited receptive fields, which is restrictive for $360{{}^{\circ}}$ FoV panoramas. The circular attention module, placed between feature extraction and cost volume construction is flexible and can be easily integrated since it maintains the input dimension. Besides, it only calculates the relations along the vertical axis, conserving more computing costs compared to global self-attention approaches. Fig. 4 (a) demonstrates our circular attention module.
In global self-attention, given an input feature map $x\in\mathbb{R}^{h\times w\times d_{in}}$ with height $h$, width $w$, and channels $d_{in}$. The output $y_{o}\in\mathbb{R}^{d_{out}}$ at position $o=(i,j)$ can be calculated as:
$y_{o}=\sum_{p\in\mathcal{N}}softmax_{p}(q_{o}^{T}k_{p})v_{p}$ (6)
where $\mathcal{N}$ is the whole location lattice, $p=(a,b)$ are all possible positions. Queries $q_{o}=W_{Q}x_{o}$, keys $k_{o}=W_{K}x_{o}$, and values $v_{o}=W_{V}x_{o}$ are all linear projections of the input $x_{o}$, where $\forall o\in\mathcal{N}$. $W_{Q},W_{K}\in\mathbb{R}^{d_{q}\times d_{in}}$, and $W_{V}\in\mathbb{R}^{d_{out}\times d_{in}}$ are all learnable weights.
However, global self-attention is extremely resource-consuming and computes $(\mathcal{O}(h^{2}w^{2}))$. Inspired by  [35, 11], we restrict the receptive field of self-attention to a local region and apply only along the vertical axis. Additionally, global self-attention doesn’t contain positional information, which is proven to be effective in many works [36, 35, 46, 54]. We incorporate positional information in the circular attention module. The output $y_{o}$ at position $o=(i,j)$ can be calculated as:
$y_{o}=\sum_{p\in\mathcal{N}_{1\times m}(o)}softmax_{p}(q_{o}^{T}k_{p}+q_{o}^{T}r_{p-o}^{q}+k_{p}^{T}r_{p-o}^{k})(v_{p}+r_{p-o}^{v})$ (7)
where $\mathcal{N}_{1\times m}(o)$ is the local $1\times m$ region centered around location $o=(i,j)$. $r^{q}_{p-o}\in\mathbb{R}^{d_{q}}$ is the learnable relative positional encoding for queries and the inner product $q^{T}_{o}r^{q}_{p-o}$ measures the compatibility from location $p$ to location $o$. Similarly, the learnable vectors $r^{k}_{p-o}\in\mathbb{R}^{d_{q}}$ and $r^{v}_{p-o}\in\mathbb{R}^{d_{out}}$ are positional encodings for keys and values. Our circular attention reduces the computation to $(\mathcal{O}(hwm))$.
For the Deep360 dataset, the feature map size after feature extraction is $h\times w\times d_{in}=256\times 128\times 32$. After a $1\times 1$ convolution is applied, the feature map is fed into a multi-head attention module, where the attention mechanism is only applied along the vertical axis. We set span $m=256$ to ensure it captures all features along the vertical axis. We use 8 heads, each producing $256\times 128\times 4$ outputs. These are concatenated to $256\times 128\times 32$, and after another $1\times 1$ convolution, the feature map is added element-wise to the original. Fig. 4 (b) illustrates how one head of the circular attention module works.
4 Experiments
4.1 Datasets
We train and evaluate our framework on Deep360 [19] and 3D60 [65], which include outdoor and indoor scenes. We evaluate both stereo matching and depth estimation. For Deep360, four $360{{}^{\circ}}$ cameras are arranged horizontally in a square. Panoramas from all four views are used for evaluation. We use six stereo pairs for training and testing. For 3D60, three $360{{}^{\circ}}$ cameras are arranged vertically in an equilateral right triangle. Panoramas from two of three views are used for evaluation. The resolutions are $1024\times 512$ and $512\times 256$, respectively.
4.2 Evaluation Metrics
Following MODE [19], we evaluate stereo matching performance using MAE (mean absolute error), RMSE (root mean square error), Px1,3,5 (percentage of outliers with pixel error >> 1, 3, 5), D1 [30] (percentage of outliers with pixel error >> 3 and >> 5%). We evaluate depth estimation performance using MAE, RMSE, AbsRel (absolute relative error), SqRel (square relative error), SILog [7](scale-invariant logarithmic error), $\delta 1,2,3$ [16] (accuracy with threshold that $\max(\tfrac{\hat{y}}{y^{\star}},\tfrac{y^{\star}}{\hat{y}})<1.25,1.25^{2},1.25^{3}$).
4.3 Implementation Details
We apply nearest-neighbor interpolation for cylindrical/cubic disparity maps generalization and bilinear interpolation for cylindrical/cubic panoramas generalization, both derived from spherical inputs.
In the stereo matching stage, cylindrical panoramas have a $360{{}^{\circ}}$ vertical FoV and a horizontal FoV of less than $180{{}^{\circ}}$. We evaluate the central part of disparity maps in the Cassini domain with horizontal FoV = $2\arctan(\pi/2)\approx 105{{}^{\circ}}$ for both datasets. This FoV yields cylindrical panorama size equivalent to spherical. In the fusion stage, we evaluate the entire omnidirectional depth map with a $360{{}^{\circ}}$ horizontal and $180{{}^{\circ}}$ vertical FoV.
4.4 Experimental Results
Training on Perspective Images and Testing on Panoramas The pre-trained models of PSMNet [2] and IGEV-Stereo [56] are trained on Scene Flow [27], which contain only perspective images. CREStereo [17], trained on mixed datasets, exhibits better generalization. The performance of stereo matching with different projections on Deep360 is shown in Tab. 2. Acquiring panoramas and their depth ground truth is difficult, the experimental results demonstrate the potential to apply stereo-matching models trained on perspective images to cylindrical panoramas.
Table 2: Quantitative results of stereo matching models pre-trained on perspective datasets evaluated on the Deep360 test dataset under different projections.
Method Projection MAE Px1 (%) D1 (%)
PSMNet [2] Cassini 2.7667 42.7912 12.6288
Cylindrical 2.6118 34.4403 10.8204
IGEV-Stereo [56] Cassini 6.5155 61.0948 29.7265
Cylindrical 4.0194 53.3429 22.8117
CREStereo [17] Cassini 4.6836 43.5014 18.5130
Cylindrical 2.1241 22.6015 11.2502
Figure 5: Depth estimation results on the Deep360 test dataset.
Comparisons with State-of-the-Art Methods We first evaluate our method against leading stereo matching networks such as PSMNet [2], AANet [57], and 360SD-Net [47], which is designed for $360{{}^{\circ}}$ stereo. We train these models on the Deep360 training dataset from scratch and test them on the Deep360 test dataset following the default experimental settings. Tab. 3 shows that our method achieves state-of-the-art performance.
For omnidirectional depth estimation, we compare our method with other multi-view omnidirectional depth estimation methods including UNiFuse [14], CSDNet [18], 360SD-Net [47], OmniMVS [52], and MODE [19]. We report the results from MODE. Tab. 4 shows that our method achieves an 18.8% MAE reduction on Deep360 and 19.9% on 3D60 compared to the previous best results, confirming its effectiveness for diverse real-world panoramas.
Table 3: Quantitative results of stereo matching methods on Deep360 and 3D60 test datasets. The top three results for each metric are highlighted with a first, second, and third background, respectively.
Dataset Method Projection Kernel Type MAE $\downarrow$ RMSE $\downarrow$ Px1 (%) $\downarrow$ Px3 (%) $\downarrow$ Px5 (%) $\downarrow$ D1 (%) $\downarrow$
Deep360 AANet [57] Cassini Regular 0.3427 1.5703 5.2050 2.1515 1.2847 1.9817
360SD-Net [47] Cassini Regular 0.5262 1.6459 3.8794 1.3389 0.8425 1.2989
PSMNet [2] Cassini Regular 0.2703 1.4790 3.3556 1.1979 0.7538 1.1708
MODE [19] Cassini Spherical 0.2309 1.4014 2.8801 1.0488 0.6562 1.0326
Ours Cylindrical Regular 0.2112 1.3903 2.5713 1.0009 0.6376 0.9828
3D60 MODE [19] Cassini Spherical 0.2258 0.5265 2.9441 0.6482 0.2978 0.6478
Ours Cylindrical Regular 0.1773 0.4654 2.2298 0.5282 0.2564 0.5279
Table 4: Quantitative results of omnidirectional depth estimation methods on Deep360 and 3D60 test datasets.
Dataset Method Kernel Type MAE $\downarrow$ RMSE $\downarrow$ AbsRel $\downarrow$ SqRel $\downarrow$ SILog $\downarrow$ $\delta 1\%\uparrow$ $\delta 2\%\uparrow$ $\delta 3\%\uparrow$
Deep360 OmniMVS [52] Regular 8.8865 59.3043 0.1073 2.9071 0.2434 94.9611 97.5495 98.2851
360SD-Net [47] Regular 11.2643 66.5789 0.0609 0.5973 0.2438 94.8594 97.2050 98.1038
CSDNet [18] Spherical 6.6548 36.5526 0.1553 1.7898 0.2475 86.0836 95.1589 97.7562
UniFuse [14] Regular 3.9193 28.8475 0.0546 0.3125 0.1508 96.0269 98.2679 98.9909
MODE [19] Spherical 3.2483 24.9391 0.0365 0.0789 0.1104 97.9636 99.0987 99.4683
Ours+Cubic Regular 5.0309 36.1907 0.0785 0.4410 0.1781 94.5960 98.1782 98.9406
Ours Regular 2.6384 21.6692 0.0304 0.1153 0.1033 98.2557 99.2101 99.5227
3D60 360SD-Net [47] Regular 0.0762 0.2639 0.0300 0.0117 1.4578 97.6751 98.6603 99.0417
CSDNet [18] Spherical 0.2067 0.4225 0.0908 0.0241 0.1273 91.9537 98.3936 99.5109
UniFuse [14] Regular 0.1868 0.3947 0.0799 0.0246 0.1126 93.2860 98.4839 99.4828
MODE [19] Spherical 0.0713 0.2631 0.0224 0.0031 0.0512 99.1283 99.7847 99.9250
Ours Regular 0.0571 0.1903 0.0199 0.0027 0.0401 99.3933 99.8506 99.9418
Fig. 5 illustrates the superior performance on Deep360, effectively handling severe distortions while preserving finer object details and the edges between the foreground and background better than MODE.
Figure 6: Depth estimation results in real-world scenarios.
Performance on Real Scenarios We evaluate our models on real-world fisheye image pairs. We reproject the fisheye images in Cassini projection, which have a vertical FoV of $189{{}^{\circ}}$ and a horizontal FoV of $120^{\circ}$. The lens projection is equidistant, and we use OpenCV with checkerboards to calibrate the camera parameters and the relative pose of the two cameras.  Fig. 6 shows the qualitative results of our models compared to MODE. For both indoor and outdoor scenes, with models trained on 3D60 and Deep360 respectively, MCPDepth demonstrates noticeable improvements over MODE, particularly in the highly distorted areas.
4.5 Ablation Study
Panorama Projection Tab. 5 demonstrates that cylindrical projection significantly outperforms spherical projection in stereo matching, even when applying spherical convolutions on spherical panoramas (MODE). Furthermore, we compare the performance of different projections on depth estimation in Tab. 4, demonstrating that cylindrical projection is the most suitable projection for regular convolutions, making it more effective for panorama stereo matching and depth estimation. These benefits may extend to other panorama-related vision tasks.
Circular Attention Tab. 6 shows that, although designed to mitigate vertical distortion for cylindrical projection, our circular attention module consistently improves performance across various panoramic projections and stereo-matching networks. This lightweight module delivers significant accuracy gains with minimal additional computation, evaluated in the Cassini domain for spherical/cylindrical panoramas and the cubic domain for cubic panoramas. For IGEV-Stereo [56], applying circular attention to the largest feature map (first scale) yields substantial performance improvements.
Table 5: Ablation study for different projections on the Deep360 test dataset. The metrics refer to disparity errors.
Method Projection MAE Px1 (%) D1 (%)
MODE [19] Cassini 0.2309 2.8801 1.0326
PSMNet [2] Cassini 0.2703 3.3556 1.1708
Cylindrical 0.2179 2.6489 1.0236
IGEV-Stereo [56] Cassini 0.3905 6.1733 1.8843
Cylindrical 0.3278 4.7958 1.7276
Table 6: Ablation study for circular attention module on the Deep360 test dataset. ”CA” denotes circular attention. The metrics refer to disparity errors.
Method Projection CA MAE Px1 (%) D1 (%)
MODE [19] Cassini 0.2309 2.8801 1.0326
Cassini 0.2210 2.7537 0.9881
PSMNet [2] Cubic 0.4471 5.0001 1.7623
Cubic 0.4196 4.6699 1.6464
PSMNet [2] Cylindrical 0.2179 2.6489 1.0236
Cylindrical 0.2112 2.5713 0.9828
IGEV-Stereo [56] Cylindrical 0.3278 4.7958 1.7276
Cylindrical 0.2265 2.9581 1.1052
5 Conclusion
We present MCPDepth, a novel two-stage framework for omnidirectional depth estimation through stereo matching from multiple cylindrical panoramas. Our comprehensive theoretical and experimental comparisons on different panoramic projections highlight the distinct advantages of cylindrical projection. Cylindrical projection maintains the linear epipolar constraint and preserves the definition of disparity as in perspective images. It effectively reduces distortion, enabling the application of stereo-matching models trained on perspective images to cylindrical panoramas. Additionally, cylindrical projection eliminates the need for customized kernels, simplifying deployment on embedded devices. Our circular attention module addresses vertical-axis distortions in cylindrical panoramas and captures $360{{}^{\circ}}$ features, and can be extended to other projections. Experimental results demonstrate that MCPDepth achieves state-of-the-art performance on both the outdoor dataset Deep360 and the indoor dataset 3D60.
Limitations Cylindrical panoramas are limited in their ability to capture a full $180{{}^{\circ}}$ horizontal FoV. As a result, at least 3 cameras are required to ensure complete coverage. Future work should explore optimizing the horizontal FoV of cylindrical panoramas to balance performance and computational resources.
Acknowledgments
We gratefully acknowledge the advanced computational resources provided by Engineering IT and Research Infrastructure Services at Washington University in St. Louis.